OpenStudy (anonymous):

Given an arithmetic sequence with a7=19 and a12=54, find a15.

OpenStudy (amistre64):

12-7 = 5 steps 54 = 19 + 5(d) ; solve for d

OpenStudy (anonymous):

7?

OpenStudy (amistre64):

54-19 = 5(d) .... 7 is good :)

OpenStudy (amistre64):

now lets use that to determine 8 steps from a7 a{15} = 19 + 8(7) i believe

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (amistre64):

youre welcome :) always dbl chk me work tho... im old and forgetful

hero (hero):

Looks right to me